Given the utility function:

U = X 3/4 . Y1/4

Estimate the demand functions of commodity X and commodity Y using Lagrange method, if it is given that price of X is Px and price of Y is Py and Income is M.


1
Expert's answer
2020-08-14T16:48:31-0400

"L=X*3/4*Y*1/4+\lambda(M-PxX-PyY)"

"\frac{\delta L}{\delta X}=3/16*Y-\lambda Px=0"

"\frac{\delta L}{\delta Y}=3/16*X-\lambda Py=0"

"\frac{\delta L}{\delta \lambda}=M-PxX-PyY=0"

"\frac{3/16*Y}{\lambda}=Px"

"\frac{3/16*X}{\lambda}=Py"

"M-\frac{3/16*Y}{\lambda}X-\frac{3/16*X}{\lambda}Y=0"

"M-\frac{3/8*YX}{\lambda}=0"

"\lambda*M=3/8*Y*X"

Commodities X and Y are substitutes does not matter their ratio you will get maximum utility